OpenStudy (hannahc234):

Will fan and medal whoever checks my work. 3 ^-2 ( ----- ) 5y^4 A. 6 ----- 10y^16 B. 10y^16 ----- 6 C. 9 ----- 25y^8 D. 25y^8 ----- 9 I think it is C.

OpenStudy (faiqraees):

wrong notice the exponent have a negative sign

OpenStudy (faiqraees):

\[\large\rm a^{-1}= \frac{1}{a}\]

OpenStudy (faiqraees):

first solve the problem as if there was no negative sign. After that just make its reciprocal.

OpenStudy (hannahc234):

I still don't get it. I understood the first one because it didn't have a negative. Negatives get me SO confused.

OpenStudy (faiqraees):

Okay should I walk you through the solution?

OpenStudy (hannahc234):

I think that would help yeah lol

OpenStudy (faiqraees):

\[\large\rm \frac{3}{5y^4}^{-2}\]\[\large\rm 3^{-2}*\frac{1}{5y^4}^{-2}\]\[\large\rm 9^{-1}*\frac{1}{25y^8}^{-1}\] So far everything clear?

OpenStudy (hannahc234):

Yes

OpenStudy (faiqraees):

\[\large\rm a^{-1}=\frac{1}{a}\]\[\large\rm \frac{1}{a}^{-1}=a\] Agree?

OpenStudy (hannahc234):

Yes

OpenStudy (faiqraees):

\[\large\rm 9^{-1}=\frac{1}{9} \] \[\large\rm \frac{1}{25y^8}^{-1}=25y^8\] Can you work out the answer now

OpenStudy (hannahc234):

So it's D?

OpenStudy (phi):

yes, it's D. The fast way is if you see a negative exponent, "flip" the fraction, and at the same time make the exponent positive so when you see \[ \left(\frac{3}{5y^4}\right)^{-2} \] flip the fraction and make the -2 a +2 \[ \left(\frac{5y^4}{3}\right)^{2} \] now do the normal thing

OpenStudy (faiqraees):

Yep it is D